# Bilinear relational structure fixes reversal curse and enables consistent model editing

### v2

@@ 4.3 Experiment 3: Relational Algebra Tests
Figure 4: Performance on relational algebra tasks. Top Row (Composition): Accuracy of inferring a composed relation using the product of the corresponding bilinear matrices (e.g., $M_{\text{husband}}\cdot M_{\text{mother}}$ to probe for ‘father’). Bottom Row (Transpose): Accuracy of inferring an inverse relation using the transpose of a bilinear matrix (e.g., $M_{\text{husband}}^{\top}$ to probe for ‘wife’).
Having established a strong bilinear signal in Experiment 2, we test whether the trained relation matrices $M_{r}$ obey basic algebraic laws that enable inverse and multi-hop reasoning. Using the matrices estimated by the bilinear probe in each layer, we evaluate two relational reasoning operations with the same scoring function $f_{r}$ and evaluation protocol as the bilinear probe: (1) Transpose/Inversion: does $M_{r}^{\top}$ act like the inverse relation $r^{-1}$? (2) Composition: does the product $M_{r_{2}}M_{r_{1}}$ act like the composed relation $r_{2}\!\circ\!r_{1}$?
@@ 4.4 Experiment 4: Model editing and its link to bilinear structure
We evaluated edited models on three metrics (Figure 5a): (1) edit success—whether the model correctly predicts (A, husband, B’); (2) logical generalization—the average success rate on entailed facts: (B’, wife, A), (C/D, father, B’), and (B’, daughter/son, C/D); and (3) locality—whether unrelated facts like (C, brother, D) remain unchanged.
@@ Results.
Figure 5b shows that both “Reversal Cursed” and “Not Reversal Cursed” models successfully learn the direct edit with 100% accuracy. However, their generalization abilities diverge dramatically: models without the reversal curse achieve high logical generalization, while those with the curse fail completely. Additionally, “Not Reversal Cursed” models maintain better locality than their counterparts.
To quantify the relationship between internal structure and editing performance, we correlate each model’s best bilinear probe accuracy (across all layers from Experiment 2) with its best logical generalization after editing (across all layers). Figure 5c reveals a strong positive correlation ($R^{2}=0.939$), demonstrating that a well-structured bilinear representation predicts a successful logical propagation of a single edit.
@@
Figure 5: Model editing generalization and its link to bilinear structure. (a) Schematic of the editing task. The fact (A, husband, B) is edited to (A, husband, B’). A successful logical generalization updates the inverse (B’, wife, A) and neighborhood relations (C/D, father, B’); (B’, daughter/son, C/D). (b) Performance after editing the target layer: Edit Success (direct change), Logical Generalization (propagation to entailed facts), and Locality (impact on unrelated facts). (c) A strong correlation ($R^{2}=0.939$) exists between a model’s best bilinear accuracy and its best logical generalization after editing. (d) Layer-wise performance of bilinear probing and logical generalization for “Not Reversal Cursed” models.
@@ 5 Discussion
Our findings establish a clear mechanistic link between the relational structure in the training data, the emergent representational geometry, and logical generalization. This has significant implications for how we understand, build, and interact with LMs.
@@
This work suggests that the perceived gap between a LM’s ability to memorize and its ability to reason may be a function of its internal knowledge structure. Phenomena like the reversal curse appear as symptoms of an LM optimized for exploiting statistical shortcuts—such as relying on the co-occurrence of terms in training data—rather than developing the logical understanding such as latent multi-hop reasoning (Yang et al., 2024a; b; Balesni et al., 2025). Our demonstration that a structured knowledge dataset with appropriate regularization can induce a transition to an algebraic, bilinear structure implies that transformer-based LMs are capable of learning more than just directional associations. This raises a crucial question for the field: Are we explicitly training models to reason, or are we hoping reasoning emerges as a side effect of scaling? Our results suggest that actively guiding the formation of structured representations, perhaps through curriculum learning, contrastive objectives, or integration with knowledge graphs during pretraining, could be a more direct path to building genuinely logical LMs.
Model Editing: Is the Model “Ready” to be Edited?
Our results reframe the challenge of model editing. Much of the current research focuses on developing more sophisticated editing algorithms, treating the model as a static object to be operated upon. We show that the success of any edit is fundamentally constrained by the pre-existing representational geometry. An editor cannot force a logically entailed update if the model lacks the necessary algebraic structure to represent that entailment. This suggests a paradigm shift: before editing, we might first need to assess whether a model is “editable” in a logically consistent way. This leads to a two-pronged approach for future research: 1) developing editing algorithms that are aware of and can leverage the model’s internal geometry, and 2) “preparing” models for editing by endowing them with structured knowledge representations during pre-training or fine-tuning.
The Double-Edged Sword of Structured Knowledge.
While a well-structured bilinear representation enables the desirable propagation of edits, it also presents a safety challenge. In a model with this structure, a single edit could trigger a cascade of unintended changes throughout its knowledge base, a phenomenon we might call “catastrophic generalization.” For example, consider a model trained on data before 2025. Updating it with a single edit changing “The current US president is Joe Biden” to “The current US president is Donald Trump” could be disastrous if the model’s generalization is too broad. It may incorrectly rewrite dozens of related facts from 2020 to 2024, presenting a distorted and false view of the world to the user. Ensuring safe model editing will therefore require promoting logical generalization as well as developing mechanisms to trace, constrain, and verify the downstream impact of any modification within this tightly coupled algebraic system.
@@
Our primary limitation is its use of LMs trained from scratch on a clean, synthetic dataset. This raises the crucial question of whether these findings scale to large, pre-trained LMs and the noisy, complex knowledge they contain. Whether similar bilinear structures exist in industrial-scale pre-trained LMs remains an open question; it is unlikely that all knowledge is encoded via a single, uniform geometry. Instead, different domains of knowledge may adopt different relational structures in their representations. Our work is a proof of concept, demonstrating that LMs are capable of forming this algebraically robust structure, although we have not verified its prevalence in existing large-scale LMs. A critical direction for future work, therefore, is to develop methodologies to diagnose how a pre-trained model decodes a specific piece of information. Such a diagnostic capability would be transformative, enabling a new paradigm of structure-aware model editing. By first identifying a fact’s local representational geometry, we could then select or design editing techniques that respect and leverage that structure, moving the field from a trial-and-error process to a more principled, geometrically informed science of knowledge modification.
@@ 6 Conclusion
We demonstrated that the reversal curse and failures in model editing generalization are not inherent limitations of LMs, but rather symptoms of an unstructured internal knowledge representation. By training transformers on a synthetic knowledge graph with appropriate regularization, we showed that they can learn a robust bilinear structure for relational knowledge. Probing experiments confirmed that this structure emerges in the middle layers and is algebraically sound, supporting relational inversion and composition. Crucially, we established a strong link between the presence of this bilinear representation and the model’s ability to both overcome the reversal curse and perform logically consistent model editing. When a fact was edited, models with this structure successfully propagated the change to entailed facts, whereas models without it failed to generalize. Our findings highlight that the path toward more reliable and editable LMs lies not just in better algorithms but in shaping the fundamental geometry of their learned knowledge representations.
@@
We include full details about our model architecture, training setup, and hyperparameter sweeps in Appendix A. Our synthetic dataset construction are described in Appendix B. We ran all experiments on a workstation with 4 A100 GPUs. Our implementation uses the GPT-NeoX library (Andonian et al., 2023) via HuggingFace Transformers Wolf et al. (2019), which is implemented in PyTorch (Paszke et al., 2019).
@@ Appendix C SVD-based Update of Relation Matrices in RESCAL
$n=1250$ (max entities; 125 families, 10 member for each family) and $d=896$ (embedding dimension) in our experiments, making the SVD-based update feasible. $M_{r}$ is obtrained from train set $A$ then evaluated on test set $B$. $B$ and $A$ have disjoint entity sets.
@@ D.2.1 Linear relation embedding
Figure 7 shows the probe’s accuracy as a function of the number of training samples ($n=10$ vs. $n=100$). The main text reports results for $n=10$ due to the high computational cost of Jacobian calculations. Here, we show that increasing the number of samples to $n=100$ does not improve performance, confirming that the poor accuracy of the linear relational embedding probe is not due to insufficient sampling.
@@
Figure 7: Linear relational embedding accuracy for $n=100$ and $n=10$ where $n$ is the number of training samples for probing.
Figure 8: Visualization of linear relational embedding probing results for each relation $r$ in with spaghetti plot ($n=100$ and $\beta=5$).
D.2.2 Tranlsational
@@ D.2.3 Bilinear
Figure 10 shows the per-relation accuracy of blinear probing. The “Not Reversal Cursed” models achieve high accuracy across all relations, peaking at mid-to-late layers. In contrast, “Reversal Cursed” models show near-zero accuracy for all relations. This indicates that learning a bilinear relational structure is strongly associated with overcoming the reversal curse and generalizes well across different relation types.

### v3

@@ 4.3 Experiment 3: Relational Algebra Tests
Figure 4: Performance on formal algebraic tests. Top Row (Composition): Accuracy of inferring a composed relation using the product of the corresponding bilinear matrices (e.g., $M_{\text{husband}}\cdot M_{\text{mother}}$ to probe for ‘father’). Bottom Row (Transpose or Inversion): Accuracy of inferring an inverse relation using the transpose of a bilinear matrix (e.g., $M_{\text{husband}}^{\top}$ to probe for ‘wife’).
Having established a strong bilinear signal in Experiment 2, we test whether the trained relation matrices $M_{r}$ obey basic algebraic laws that enable inverse and multi-hop reasoning. Using the matrices estimated by the bilinear probe in each layer, we evaluate two relational reasoning operations with the same scoring function $f_{r}$ and evaluation protocol as the bilinear probe: (1) Transpose or Inversion: does $M_{r}^{\top}$ act like the inverse relation $r^{-1}$? (2) Composition: does the product $M_{r_{2}}M_{r_{1}}$ act like the composed relation $r_{2}\!\circ\!r_{1}$?
@@ 4.4 Experiment 4: Model editing and its link to bilinear structure
We evaluated edited models on three metrics (Figure 5a): (1) edit success, which measures whether the model correctly predicts the primary update (A, husband, B’); (2) logical generalization, defined as the average success rate on the entailed facts: (B’, wife, A), (C/D, father, B’), and (B’, daughter/son, C/D); and (3) locality, which assesses whether unrelated facts like (C, brother, D) remain unchanged.
@@ Results.
As illustrated in Figure 5b, both “Reversal Cursed” and “Not Reversal Cursed” models achieve near-perfect accuracy on the primary edit. However, their capacities for logical propagation diverge sharply: models without the reversal curse exhibit robust generalization to entailed facts, whereas reversal-cursed models fail almost entirely. Furthermore, “Not Reversal Cursed” models demonstrate superior locality, preserving the integrity of unrelated factual knowledge more effectively than their counterparts.
To quantify the relationship between internal structure and editing performance, we correlate each model’s best bilinear probe accuracy (across all layers from Experiment 2) with its best logical generalization after editing (across all layers). Figure 5c reveals a robust positive correlation ($R^{2}=0.939$), establishing that a well-structured bilinear representation predicts a successful logical propagation of a single edit.
Figure 5: Editing generalization and its link to bilinear structure. (a) Schematic of the editing task. The primary fact (A, husband, B) is edited to (A, husband, B’). Successful logical generalization requires updating the inverse (B’, wife, A) and neighborhood relations (C/D, father, B’); (B’, daughter/son, C/D). (b) Post-edit performance metrics: Edit Success (direct change), Logical Generalization (propagation to entailed facts), and Locality (impact on unrelated facts). (c) Correlation between a model’s peak bilinear probe accuracy and its maximum logical generalization performance ($R^{2}=0.939$). (d) Layer-wise performance of bilinear probing and logical generalization for “Not Reversal Cursed” models.
@@ 5 Discussion
Our findings establish a clear mechanistic link between relational structures in training data, emergent representational geometry, and logical generalization. These results have significant implications for how we understand, build, and control language models.
Model Editing: Is the Model “Ready” to be Edited?
Our results reframe the fundamental challenge of model editing. Much of the current research focuses on developing more sophisticated editing algorithms, treating the model as a static object to be operated upon. We show that the success of any edit is fundamentally constrained by the model’s pre-existing representational geometry: an editor cannot force a logically entailed update if the underlying architecture lacks the necessary algebraic structure required to represent that entailment.
However, the existence of structure alone is not enough; the editing method must also preserve this structural integrity. While we show that fine-tuning can successfully propagate edits when a bilinear structure is present, recent evidence suggests that aggressive editing methods can “shatter” the underlying graph topology (Nishi et al., 2025). Thus, while bilinear structure is a key indicator of logical generalization, realizing this potential requires editing algorithms that respect—rather than destroy—the model’s internal algebraic integrity. This suggests a paradigm shift: before editing, one must first assess whether a model is “editable” in a logically consistent way. This leads to a two-pronged approach for future research: 1) the development of geometry-aware editing algorithms that leverage internal relational structures, and 2) “architectural preparation” strategies that explicitly endow models with structured knowledge representations during pre-training or fine-tuning.
@@
This work suggests that the perceived gap between a model’s ability to memorize and its ability to reason may be a function of its internal knowledge structure. Phenomena like the reversal curse can be viewed as consequences of optimizing models to exploit statistical shortcuts (e.g., term co-occurrence in the training data) rather than to develop capabilities that support logical understanding, such as latent multi-hop reasoning (Yang et al., 2024; 2025a; Balesni et al., 2025). Our demonstration that a structured knowledge dataset with appropriate regularization can induce a transition to an algebraic, bilinear structure implies that Transformers are capable of learning more than just directional associations. This raises a crucial question for the field: Are we explicitly training models to reason, or are we hoping reasoning emerges as a side effect of scaling? Our results suggest that actively guiding the formation of structured representations, perhaps through curriculum learning, contrastive objectives, or integration with knowledge graphs during pretraining (Yang et al., 2025b), could be a more direct path to building genuinely logical AI.
Mechanistic Interpretation: The Attention Head Hypothesis.
While our analysis relies on readout probes and algebraic validation, we hypothesize that the “bilinear structure” we detect corresponds to interaction patterns implemented within the self-attention layers. Because attention computes interactions between pairs of tokens through multiplicative comparisons, it is inherently bilinear (e.g., terms of the form $x^{\top}Ay$). It is therefore plausible that particular attention heads are responsible for encoding these relational structures. Recently, Elhelo & Geva (2025) have shown that specific attention heads encode relational lookups and tend to appear in mid to late layers, aligning with our observation that bilinear structure emerges in those layers. This suggests that the absence of such specialized attention heads may be a bottleneck in reversal-cursed models. Future work could test this hypothesis by intervening on attention heads whose behavior aligns with our estimated $M_{r}$ matrices.
@@
Our primary limitation is the use of language models trained from scratch on a clean, synthetic dataset. This raises the crucial question of whether these findings scale to large, pre-trained language models and the noisy, complex knowledge they contain. Whether similar bilinear structures exist in industrial-scale pre-trained language models remains an open question; it is unlikely that all knowledge is encoded via a single, uniform geometry. Instead, different domains of knowledge may adopt different relational structures in their representations. Our work is a proof of concept, demonstrating that language models are capable of forming this algebraically robust structure, although we have not verified its prevalence in existing large-scale models. A critical direction for future work, therefore, is to develop methodologies to diagnose how a pre-trained model decodes a specific piece of information. Such a diagnostic capability would be transformative, enabling a new paradigm of structure-aware model editing. By first identifying a fact’s local representational geometry, we could then select or design editing techniques that respect and leverage that structure, moving the field from a trial-and-error process to a more principled, geometrically informed science of knowledge modification.
@@ 6 Conclusion
We demonstrated that the reversal curse and failures in model editing generalization are not inherent limitations of Transformers, but rather symptoms of an unstructured internal knowledge representation. By training Transformers on a synthetic knowledge graph with appropriate regularization, we showed that they can learn a robust bilinear structure for relational knowledge. Probing experiments confirmed that this structure emerges in the middle layers and is algebraically sound, supporting relational inversion and composition. Crucially, we established a strong link between the presence of this bilinear representation and the model’s ability to both overcome the reversal curse and perform logically consistent model editing. When a fact was edited, models with this structure successfully propagated the change to entailed facts, whereas models without it failed to generalize. Our findings highlight that the path toward more reliable and editable language models lies not just in better algorithms but in shaping the fundamental geometry of their learned knowledge representations.
@@
We include complete details about our model architecture, training setup, and hyperparameter sweeps in Appendix A. Our synthetic dataset construction is described in Appendix B. We ran all experiments on a workstation with 4 A100 GPUs. Our implementation uses the GPT-NeoX library (Andonian et al., 2023) via HuggingFace Transformers Wolf et al. (2019), which is implemented in PyTorch (Paszke et al., 2019).
@@ Appendix C SVD-based Update of Relation Matrices in RESCAL
$n=1250$ (max entities; 125 families, 10 members per family) and $d=896$ (embedding dimension) in our experiments, making the SVD-based update feasible. $M_{r}$ is obtained from train set $A$ then evaluated on test set $B$. $B$ and $A$ have disjoint entity sets.
@@ D.2.1 Linear relation embedding
Figure 7 shows the probe’s accuracy as a function of the number of training samples per relation, $n$ ($n=10$, $n=100$, and $n=500$). The main text reports results for $n=10$ due to the high computational cost of Jacobian calculations. Here, we show that increasing the number of samples up to $n=500$ does not improve performance, confirming that the poor accuracy of the linear relational embedding probe is not due to insufficient sampling. We also run our bilinear probing for $n=10$, $n=100$, and $n=500$, which shows that insufficient $n$ leads to underfitting.
@@
Figure 7: Bilinear and linear relational embedding accuracy for $n=10$, $n=100$, and $n=500$, where $n$ denotes the number of training samples per relation.
Figure 8: Visualization of linear relational embedding probing results for each relation $r$ as a spaghetti plot ($n=100$ and $\beta=5$).
D.2.2 Translational
@@ D.2.3 Bilinear
Figure 10 shows the per-relation accuracy of bilinear probing. The “Not Reversal Cursed” models achieve high accuracy across all relations, peaking at mid-to-late layers. In contrast, “Reversal Cursed” models show near-zero accuracy for all relations. This indicates that learning a bilinear relational structure is strongly associated with overcoming the reversal curse and generalizes well across different relation types.